An abe-see·nce , belenud Y has い208 1-3°. Yu, s..me phese vol*ejes
Answer #1

-Van nVab ca Cn uS consideY volHuge Souyce as - Conne show in bn, ave Yes Specti ve ine n. etween e aye Calle ase voltages .bc208-30-120 bc= 208-SO ca= 208-SO-120 Ya= 208 )-270 cse volHoyex oltgd 20 o88 -30 bn = 120.088-so 3 en120 088 1-270